# A Combination Model Based on Sequential General Variational Mode Decomposition Method for Time Series Prediction

### v1

@@ Abstract
Index Terms:
non-stationary time series, combination forecasting method, SGVMD, ARIMA, signal separation.I Introduction
@@
Financial time series forecasting is an important method for studying financial market volatility[1, 2]. By using historical data generated by financial markets to establish predictive models[3], we can explore inherent volatility patterns and predict future trends[4, 5]. Financial time series, as a special type of time series[6, 7], have commonalities with time series, such as obtaining noise that may be present in the data[8, 9]. It also has the particularity of financial data, such as the nonlinear chaotic relationship between data, making it more difficult to grasp its fluctuation patterns[10, 11]. Therefore, how to accurately reveal the trend of changes in financial time series and make reasonable predictions of financial time series is a key concern for the academic community and practitioners[12, 13].
@@
The research objects of financial time series forecasting are diverse[14, 15, 16, 17], and the research methods have also undergone extensive evolution[18], from statistical models to intelligent models, from single models to composite models. Statistical models[19], represented by the Differential Autoregressive Moving Average (ARIMA) model and the Holt-Winters forecasting method model[20, 21, 22, 23, 24], have a wide range of applications. For example, combining ARIMA with various decomposition algorithms such as Empirical Mode Decomposition (EMD) and Variational Mode Decomposition (VMD) for predicting complex time series; For example, using an improved ARMA model for stock market forecasting. However, the above models need to be built on the basis of stable sequence data, and usually require testing and preprocessing of the original data, which may lead to the loss of some hidden information, especially in big data samples, and this disadvantage is easily magnified.
@@
With the development of computer technology, intelligent models represented by artificial neural networks (ANNs) are gradually emerging. This type of model is good at handling incomplete, fuzzy, uncertain, or irregular data, and has a good fit to nonlinear relationships. Shallow neural networks represented by backpropagation neural networks (BPNN) and shallow machine learning represented by support vector machines (SVM) are also widely used in financial market prediction. However, shallow neural networks do not consider the temporal nature of data, and financial time series often have certain long-term dependencies. Therefore, recurrent neural networks (RNNs) with memory function have become the latest choice. The output of RNN at a certain moment can be used as input to feedback to neurons again, and this cascade structure is very suitable for time series data, which can preserve the dependency relationships in the data. However, in a large amount of practice, it has been found that standard RNNs are difficult to achieve long-term information preservation, and may be plagued by vanishing or exploding gradients during training. Therefore, the Long Short Term Memory Neural Network (LSTM) used to improve the standard RNN is more efficient in practical applications of financial time series prediction[25].
@@
Regardless of the individual model, it has its own advantages and limitations, so we will propose an improved VMD model. The formation of this method can be divided into two steps: (1) introducing the generalized form of the Variational Mode Decomposition (GVMD) ; (2) proposing the Sequentialized and generalized Variational Mode Decomposition (SGVMD) for the case of unknown modal number[26]. These two algorithms are crucial for time series decomposition. In the decomposition step of combination model, the extracted components are predicted. For the trend component, an ARIMA model is directly constructed for prediction. For the narrowband FM-AM component , the upper and lower envelope lines are first extracted, and then separate ARIMA models are constructed for each envelope line for prediction[27, 28]. The operation of time series that we propose, called SGVMD-E-ARIMA model because of the envelop prediction process among[29, 30]. We also could predict every component directly, called SGVMD-ARIMA model. Decomposition models can effectively explore the implicit linear relationships in data, while adding constraints has its own uniqueness in handling nonlinear relationships. However, the large amount of data generated in financial markets is difficult to determine whether it is a completely linear or nonlinear relationship, so it is necessary to combine the advantages of decomposition models and prediction models to construct a combination model[31, 32]. The common approach to constructing composite models is to decompose the data, fit the linear and nonlinear parts separately through a constrained decomposition model and a prediction model, and then superimpose them to obtain the prediction results. Based on this idea, we built a combination model to predict the financial time series, and confirmed that the combination model has advantages over a single model[33, 34].
@@
The above research has laid the foundation for the selection of research content and methods in this article. However, the existing research mostly focuses on a certain financial time series, such as sales, stock index, exchange rate, etc. It is difficult to prove whether the same model is universal for different financial time series[35]. In addition, existing combination models are often combined in a linear manner[36, 37], that is, the decomposition model part and the prediction model part are directly coupled. But if there is no direct coupling relationship between the two parts, this combination method is likely to lead to a decrease in predictive performance[38]. Therefore, this article constructs a decomposition combination prediction method with added constraints, namely the SGVMD-ARIMA model, aiming to combine the advantages of the two models to predict various financial time series such as sales, and study the prediction accuracy and universality of this combination model in different financial time series[39, 40].
II Methodology
II-A Variational Mode Decomposition (VMD)
@@
The Variational Mode Decomposition (VMD) algorithm, a powerful signal processing technique, has gained prominence in extracting inherent modes from complex signals[41, 42]. VMD employs a variational optimization framework to decompose a signal into a set of adaptive modes characterized by their respective center frequencies and amplitudes[43]. The method is particularly adept at handling non-stationary and nonlinear signals, making it valuable in various applications, such as biomedical signal processing and communication systems[44, 45].
@@
II-B General Variational Mode Decomposition (GVMD) and Sequential General Variational Mode Decomposition (SGVMD)
@@ II-C ARIMA model
II-D Sequential General Variational Mode Decomposition Method for Time Series components prediction
Generally, the components extracted by SGVMD can be divided into trend components and AM-FM components. This prediction method is suitable for predicting the AM-FM components extracted by SGVMD, and the trend component can be directly predicted using ARIMA. The implementation process of the SGVMD-E-ARIMA is introduced as follows: (1) Extract the upper and lower envelope of each component after the decomposition by the SGVMD algorithm. Here, the Hilbert transform and local interpolation method will be used, and the result is shown in Fig. 1:
Fig. 1: Modal components and their upper and lower envelope lines.
(2) For each extracted envelope, build a corresponding ARIMA model for prediction, as shown in Fig. 2:
Fig. 2: Prediction results of each envelope line.
(3) After obtaining the predicted results of the upper and lower envelopes, the prediction of the AM-FM component can be obtained by combining with the frequency-modulated (FM) component, as shown in Fig. 3:
Fig. 3: Prediction results of AM-FM components.
We propose the combination prediction method that uses SGVMD to decompose a time series, and then operate every component by predicting the AM-FM component. This process is called the SGVMD-ARIMA. In the decomposition phase, SGVMD regards the time series as a non-stationary mixed signal and uses SGVMD to decompose it into trend signals and AM-FM signals. In the prediction phase, this method distinguishes between the two types of components. For the trend signal, the ARIMA method is directly applied, while for the AM-FM component, the amplitude envelope is obtained, the envelope is predicted, and then combined with the frequency modulation component to obtain more accurate results. This process is called the SGVMD-E-ARIMA. The meaning of E in it refers to the envelope prediction process.
These two processes are given by Fig. 4 beneath.
Fig. 4: SGVMD-ARIMA and SGVMD-E-ARIMA Process
III Experimental results and evaluation
@@
III-A historical sales of an online store
The selected time series in this example is the historical sales of a certain product in an online store, and a segment of the time series that can highlight the trend is selected as the experimental dataset. The total number of data points is 1000, and its time series plot is shown in Fig. 5:
Fig. 5: Time series plot of experimental data for Example A.
The data with a length of 1000 is divided into a training set and a test set, with the ratio of 9:1. For the single ARIMA model method, different parameter combinations are compared to obtain the best fitting effect of the ARIMA(9,1,8) model on the original sequence. Since the seasonal variation of the time series in this example is proportional to the sequence level, the Holt-Winter method adopts a multiplicative model to model the sequence. After multiple experiments, the optimal parameter combination for predicting this time series is obtained as =0.91, =0.5, and =1. For the LSTM, number of hidden units is 200, learn-rate-drop-period is 400, learn-rate-drop-factor is 0.15. Regarding the EMD-ARIMA method, we first decompose the original time series into 7 IMF components, including residuals, through EMD. Then, we construct corresponding ARIMA models for each of these 7 components for prediction. SGVMD-ARIMA also belongs to a combination prediction model. Firstly, the SGVMD algorithm is used to separate the various components of the original time series. Several principal components (as shown in Fig. 5) are selected, and each of these principal components is predicted using the ARIMA method. Finally, the results are combined.
@@
Firstly, the SGVMD algorithm is applied to decompose the time series. In this process, we need to adjust the penalty factors and continuously until the extracted modal components meet the previous assumptions. After the SGVMD decomposition of the original time series, four components can be obtained, as shown in Fig. 6. The result in Fig. 6 conforms to the previous assumption: component 1 is the tendency component, while the other three components are AM-FM components with certain amplitude and central frequency.
Fig. 6: The results of IMF component analysis for the Example A.
Secondly, in the envelope prediction process, for the AM-FM components, the upper and lower envelope lines are first obtained, and ARIMA models are constructed separately for prediction. Finally, the predicted envelope function is multiplied by the pure frequency modulation signal of the component to obtain the predicted sequence. In the process of restoring the predicted sequence, there is a "lag" phenomenon in the phase. Therefore, in order to optimize the prediction effect, the phase directly obtained from the spectrum needs to undergo some minor processing, such as rounding. The prediction results of each component after optimization are shown in Fig. 7. It can be seen from the figure that the optimized predicted sequence basically matches the original sequence, which also proves the feasibility of the new prediction method. For the trend component 1, the ARIMA algorithm is directly applied for prediction.
Fig. 7: Optimized prediction results for each AM-FM component.
III-A1 Prediction Results
@@
III-A2 Evaluation and Analysis of Prediction Results
@@
III-B The Production of Australian Beer
@@
Fig. 9: The time series plot of Australian beer production.
@@
Fig. 10: The results of principal IMF analysis for the Example B.
III-B1 Prediction Results
@@
III-B2 Evaluation and Analysis of Prediction Results
The following information can be inferred from the table: (1) The SGVMD-ARIMA combination prediction method is the best-performing method in terms of evaluation indicators, while the Holt-Winter method has the least desirable prediction effect. (2) Although LSTM is a machine learning algorithm, the prediction effect of a single prediction method is still inferior to that of a combination prediction model. (3) From the perspective of sequence decomposition, the SGVMD algorithm better preserves sequence information compared to the EMD algorithm. (4) From the perspective of component prediction, directly predicting the time sequence by constructing an ARIMA model is more effective. There are two reasons for this: firstly, the problem of phase "lag" was not well resolved in this instance; secondly, the prediction performance of the ARIMA model for the envelope line was not satisfactory, leading to errors in the amplitude of the predicted sequence.
IV Conclusion And Prospect
@@
Based on the research in this paper, the SGVMD-ARIMA and the SGVMD-E-ARIMA will have broad application prospects in the prediction of non-stationary time series, especially in the field of economic analysis. Because the economic sequence is easily affected by external factors such as culture and society, it is highly non-stationary and tends to show a certain trend as time progresses. However, this method still has some shortcomings, and future improvement and improvement work will mainly focus on the following aspects: (1) Setting of penalty factors. Although SGVMD algorithm considers the composition of loss function more comprehensively, it realizes the sequential separation of components. Although it has a relatively wide range of allowable values of penalty factors, it still has the problem that the setting of penalty factors depends on experience. Therefore, the de-experienced research of penalty factors is the focus of the next step. (2) Construction of characteristic items. This is the key factor for the success of SGVMD algorithm. In the new algorithm proposed in this paper, for non-stationary time series, "narrow band" is added to the loss function as a prior feature of the component. But not all time series components can meet this assumption. Different practical problems are needed to flexibly adjust feature items to balance particularity and generality. (3) Improve the problem of phase lag. In the experiment, using the new synthetic prediction method to predict the AM-FM component will cause the phase "lag". The reason is that the unavoidable multi frequency components in the calculation of component center frequency cannot be processed, so the recovered components are not pure single frequency components, but complex frequency components with a certain change trend. The center frequency used to recover the envelope is only a certain basic frequency, so the separation parameter requirements will be more stringent. The method used in this paper is relatively crude and needs to be further analyzed and improved.

### v2

@@ Abstract
keywords:
[…]
1 Introduction
@@
Financial time series forecasting is an important method for studying financial market volatility1, 2. By using historical data generated by financial markets to establish predictive models3, we can explore inherent volatility patterns and predict future trends4, 5. Financial time series, as a special type of time series6, 7, have commonalities with time series, such as obtaining noise that may be present in the data8, 9. It also has the particularity of financial data, such as the nonlinear chaotic relationship between data, making it more difficult to grasp its fluctuation patterns10, 11. Therefore, how to accurately reveal the trend of changes in financial time series and make reasonable predictions of financial time series is a key concern for the academic community and practitioners12, 13.
@@
The research objects of financial time series forecasting are diverse14, 15, 16, 17, and the research methods have also undergone extensive evolution18, from statistical models to intelligent models, from single models to composite models. Statistical models19, represented by the Differential Autoregressive Moving Average (ARIMA) model and the Holt-Winters forecasting method model20, 21, 22, 23, 24, have a wide range of applications. For example, combining ARIMA with various decomposition algorithms such as Empirical Mode Decomposition (EMD) and Variational Mode Decomposition (VMD) for predicting complex time series; For example, using an improved ARMA model for stock market forecasting. However, the above models need to be built on the basis of stable sequence data, and usually require testing and preprocessing of the original data, which may lead to the loss of some hidden information, especially in big data samples, and this disadvantage is easily magnified.
@@
With the development of computer technology, intelligent models represented by artificial neural networks (ANNs) are gradually emerging. This type of model is good at handling incomplete, fuzzy, uncertain, or irregular data, and has a good fit to nonlinear relationships. Shallow neural networks represented by backpropagation neural networks (BPNN) and shallow machine learning represented by support vector machines (SVM) are also widely used in financial market prediction. However, shallow neural networks do not consider the temporal nature of data, and financial time series often have certain long-term dependencies. Therefore, recurrent neural networks (RNNs) with memory function have become the latest choice. The output of RNN at a certain moment can be used as input to feedback to neurons again, and this cascade structure is very suitable for time series data, which can preserve the dependency relationships in the data. However, in a large amount of practice, it has been found that standard RNNs are difficult to achieve long-term information preservation, and may be plagued by vanishing or exploding gradients during training. Therefore, the Long Short Term Memory Neural Network (LSTM) used to improve the standard RNN is more efficient in practical applications of financial time series prediction25.
@@
Regardless of the individual model, it has its own advantages and limitations, so we will propose an improved VMD model. The formation of this method can be divided into two steps: (1) introducing the generalized form of the Variational Mode Decomposition (GVMD) ; (2) proposing the Sequentialized and generalized Variational Mode Decomposition (SGVMD) for the case of unknown modal number26. These two algorithms are crucial for time series decomposition. In the decomposition step of combination model, the extracted components are predicted. For the trend component, an ARIMA model is directly constructed for prediction. For the narrowband FM-AM component , the upper and lower envelope lines are first extracted, and then separate ARIMA models are constructed for each envelope line for prediction27, 28. The operation of time series that we propose, called SGVMD-E-ARIMA model because of the envelop prediction process among29, 30. We also could predict every component directly, called SGVMD-ARIMA model. Decomposition models can effectively explore the implicit linear relationships in data, while adding constraints has its own uniqueness in handling nonlinear relationships. However, the large amount of data generated in financial markets is difficult to determine whether it is a completely linear or nonlinear relationship, so it is necessary to combine the advantages of decomposition models and prediction models to construct a combination model31, 32. The common approach to constructing composite models is to decompose the data, fit the linear and nonlinear parts separately through a constrained decomposition model and a prediction model, and then superimpose them to obtain the prediction results. Based on this idea, we built a combination model to predict the financial time series, and confirmed that the combination model has advantages over a single model33, 34.
@@
The above research has laid the foundation for the selection of research content and methods in this article. However, the existing research mostly focuses on a certain financial time series, such as sales, stock index, exchange rate, etc. It is difficult to prove whether the same model is universal for different financial time series35. In addition, existing combination models are often combined in a linear manner36, 37, that is, the decomposition model part and the prediction model part are directly coupled. But if there is no direct coupling relationship between the two parts, this combination method is likely to lead to a decrease in predictive performance38. Therefore, this article constructs a decomposition combination prediction method with added constraints, namely the SGVMD-ARIMA model, aiming to combine the advantages of the two models to predict various financial time series such as sales, and study the prediction accuracy and universality of this combination model in different financial time series39, 40.
2 Methodology
2.1 Variational Mode Decomposition (VMD)
@@
The Variational Mode Decomposition (VMD) algorithm, a powerful signal processing technique, has gained prominence in extracting inherent modes from complex signals41, 43. VMD employs a variational optimization framework to decompose a signal into a set of adaptive modes characterized by their respective center frequencies and amplitudes44. The method is particularly adept at handling non-stationary and nonlinear signals, making it valuable in various applications, such as biomedical signal processing and communication systems45, 42.
@@
2.2 General Variational Mode Decomposition (GVMD) and Sequential General Variational Mode Decomposition (SGVMD)
@@ 2.3 ARIMA model
2.4 Sequential General Variational Mode Decomposition Method for Time Series components prediction
Generally, the components extracted by SGVMD can be divided into trend components and AM-FM components. This prediction method is suitable for predicting the AM-FM components extracted by SGVMD, and the trend component can be directly predicted using ARIMA. The implementation process of the SGVMD-E-ARIMA is introduced as follows:(1) Extract the upper and lower envelope of each component after the decomposition by the SGVMD algorithm. Here, the Hilbert transform and local interpolation method will be used, and the result is shown in Fig. (a);(2) For each extracted envelope, build a corresponding ARIMA model for prediction, as shown in Fig. (b);(3) After obtaining the predicted results of the upper and lower envelopes, the prediction of the AM-FM component can be obtained by combining with the frequency-modulated (FM) component, as shown in Fig. 1:
We propose the combination prediction method that uses SGVMD to decompose a time series, and then operate every component by predicting the AM-FM component. This process is called the SGVMD-ARIMA. In the decomposition phase, SGVMD regards the time series as a non-stationary mixed signal and uses SGVMD to decompose it into trend signals and AM-FM signals. In the prediction phase, this method distinguishes between the two types of components. For the trend signal, the ARIMA method is directly applied, while for the AM-FM component, the amplitude envelope is obtained, the envelope is predicted, and then combined with the frequency modulation component to obtain more accurate results. This process is called the SGVMD-E-ARIMA. The meaning of E in it refers to the envelope prediction process. These two processes are given by Fig. 2 beneath.
3 Experimental results and evaluation
@@
3.1 historical sales of an online store
The selected time series in this example is the historical sales of a certain product in an online store, and a segment of the time series that can highlight the trend is selected as the experimental dataset. The total number of data points is 1000, and its time series plot is shown in Fig. (a):
The data with a length of 1000 is divided into a training set and a test set, with the ratio of 9:1. For the single ARIMA model method, different parameter combinations are compared to obtain the best fitting effect of the ARIMA(9,1,8) model on the original sequence. Since the seasonal variation of the time series in this example is proportional to the sequence level, the Holt-Winter method adopts a multiplicative model to model the sequence. After multiple experiments, the optimal parameter combination for predicting this time series is obtained as =0.91, =0.5, and =1. For the LSTM, number of hidden units is 200, learn-rate-drop-period is 400, learn-rate-drop-factor is 0.15. Regarding the EMD-ARIMA method, we first decompose the original time series into 7 IMF components, including residuals, through EMD. Then, we construct corresponding ARIMA models for each of these 7 components for prediction. SGVMD-ARIMA also belongs to a combination prediction model. Firstly, the SGVMD algorithm is used to separate the various components of the original time series. Several principal components (as shown in Fig. (b) are selected, and each of these principal components is predicted using the ARIMA method. Finally, the results are combined.
@@
Firstly, the SGVMD algorithm is applied to decompose the time series. In this process, we need to adjust the penalty factors and continuously until the extracted modal components meet the previous assumptions. After the SGVMD decomposition of the original time series, four components can be obtained, as shown in Fig. (b). The result in Fig. (b) conforms to the previous assumption: component 1 is the tendency component, while the other three components are AM-FM components with certain amplitude and central frequency.
Secondly, in the envelope prediction process, for the AM-FM components, the upper and lower envelope lines are first obtained, and ARIMA models are constructed separately for prediction. Finally, the predicted envelope function is multiplied by the pure frequency modulation signal of the component to obtain the predicted sequence. In the process of restoring the predicted sequence, there is a ”lag” phenomenon in the phase. Therefore, in order to optimize the prediction effect, the phase directly obtained from the spectrum needs to undergo some minor processing, such as rounding. The prediction results of each component after optimization are shown in Fig. (c). It can be seen from the figure that the optimized predicted sequence basically matches the original sequence, which also proves the feasibility of the new prediction method. For the trend component 1, the ARIMA algorithm is directly applied for prediction.
3.1.1 Prediction Results
@@
3.1.2 Evaluation and Analysis of Prediction Results
@@
3.2 The Production of Australian Beer
@@
3.2.1 Prediction Results
@@
3.2.2 Evaluation and Analysis of Prediction Results
The following information can be inferred from the table: (1) The SGVMD-ARIMA combination prediction method is the best-performing method in terms of evaluation indicators, while the Holt-Winter method has the least desirable prediction effect. (2) Although LSTM is a machine learning algorithm, the prediction effect of a single prediction method is still inferior to that of a combination prediction model. (3) From the perspective of sequence decomposition, the SGVMD algorithm better preserves sequence information compared to the EMD algorithm. (4) From the perspective of component prediction, directly predicting the time sequence by constructing an ARIMA model is more effective. There are two reasons for this: firstly, the problem of phase ”lag” was not well resolved in this instance; secondly, the prediction performance of the ARIMA model for the envelope line was not satisfactory, leading to errors in the amplitude of the predicted sequence.
4 Conclusion And Prospect
@@
Based on the research in this paper, the SGVMD-ARIMA and the SGVMD-E-ARIMA will have broad application prospects in the prediction of non-stationary time series, especially in the field of economic analysis. Because the economic sequence is easily affected by external factors such as culture and society, it is highly non-stationary and tends to show a certain trend as time progresses. However, this method still has some shortcomings, and future improvement and improvement work will mainly focus on the following aspects: (1) Setting of penalty factors. Although SGVMD algorithm considers the composition of loss function more comprehensively, it realizes the sequential separation of components. Although it has a relatively wide range of allowable values of penalty factors, it still has the problem that the setting of penalty factors depends on experience. Therefore, the de-experienced research of penalty factors is the focus of the next step. (2) Construction of characteristic items. This is the key factor for the success of SGVMD algorithm. In the new algorithm proposed in this paper, for non-stationary time series, ”narrow band” is added to the loss function as a prior feature of the component. But not all time series components can meet this assumption. Different practical problems are needed to flexibly adjust feature items to balance particularity and generality. (3) Improve the problem of phase lag. In the experiment, using the new synthetic prediction method to predict the AM-FM component will cause the phase ”lag”. The reason is that the unavoidable multi frequency components in the calculation of component center frequency cannot be processed, so the recovered components are not pure single frequency components, but complex frequency components with a certain change trend. The center frequency used to recover the envelope is only a certain basic frequency, so the separation parameter requirements will be more stringent. The method used in this paper is relatively crude and needs to be further analyzed and improved.